link1234:

(5)(3-1) + 4 is i need help

XioGonz:

First you solve for 3-1 3-1=2 5*2=10 10+4=14 answer would be 14

link1234:

tysm

XioGonz:

Np

link1234:

sorry i didn't fully anwers a question

link1234:

for AnakKrakatau1883

XioGonz:

@link1234 wrote:
sorry i didn't fully anwers a question
It's all good, Just make sure to do so next time (:

link1234:

tysm:)